OpenStudy (anonymous):

.If an investment of $2,200 is made in an account that earns 2.6% interest compounded continuously, how much money will be in the account after 12 years?

OpenStudy (anonymous):

A = P(1 + rt) •A = Total Accrued Amount (principal + interest) •P = Principal Amount •I = Interest Amount •r = Rate of Interest per year in decimal; r = R/100 •t = Time Period involved in months or years A = 2200(1 + (0.026 × 12)) = 2886.4 A = $ 2,886.40

OpenStudy (anonymous):

If you don't understand that, you could always use an interest calculator online to solve it.